OpenStudy (anonymous):

A bag contains only black marbles and white marbles. Ezra reached into the bag without looking, got 30 marbles, and found that 12 of the marbles were black. Ezra is going to reach in the bag again and get one more marble. What is the experimental probability that the marble will be black?

OpenStudy (neer2890):

what is the total number of marbles in the bag?

OpenStudy (anonymous):

12/30= 2/5 = 0.4 if we don't know the total number of marbles in the bag and how many are black and how many are white.